# Non-Stationary Lipschitz Bandits

### v1

@@ D.7 Upper bounding (B)
Proof of Proposition 10.
Let $[\tau_{l,m},\tau_{l,m+1}[$ be a block, and let $B\in\mathcal{T}_{m}$ such that $s_{l,m}(B)<\tau_{l,m}+8^{m}$. Let $d_{i,j}$ the integer satisfying
@@
$\geq\frac{c_{6}}{8}\log(T)\,,$
where we use the fact that there is at least one round $t\in[s_{i,j}^{m}(B),\tilde{s}_{i,j}^{m}(B)]$ such that $t-\tau_{l,m}\equiv 0[8^{d_{ij}}]$, i.e. that at least a replay at depth $d_{i,j}$ that can be scheduled in $[s_{i,j}^{m}(B),s_{i,j+1}(B)]$, and where we used Definition 13 in the last inequality. Setting ${\color[rgb]{0.68,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.68,0,0}\pgfsys@color@cmyk@stroke{0}{0.87}{0.68}{0.32}\pgfsys@color@cmyk@fill{0}{0.87}{0.68}{0.32}c_{6}=192}$, we have
@@
Therefore, using Proposition 10, for any bin $B\in\mathcal{T}_{m}$, on event $\mathcal{E}_{1}\cap\mathcal{E}_{2}(\tau_{l,m})$, we have that there exists a bad segment $[s_{i,j_{0}}^{m}(B),s_{i,j_{0}+1}^{m}(B)]$ such that $s_{i,j_{0}+1}^{m}(B)<s_{l,m}(B)<\tau_{l,m+1}$, and there exists a $\texttt{Replay}(t_{\mathrm{start}},d_{i,j_{0}})$ starting within this segment, $t_{\mathrm{start}}\in[s_{i,j_{0}}^{m}(B),s_{i,j_{0}+1}^{m}(B)]$ and $d_{i,j_{0}}$ satisfying $8^{d_{i,j_{0}}}\leq s_{i,j_{0}+1}^{m}(B)-s_{i,j_{0}}^{m}(B)\leq 8^{d_{i,j_{0}}+2}$. By Proposition 9, this implies that bin $B$ is evicted by round $s_{i,j_{0}+1}^{m}(B)$. Therefore, on $\mathcal{E}_{1}\cap\mathcal{E}_{2}(\tau_{l,m})$, using the definition of the bad round $s_{l,m}(B)<\tau_{l,m+1}$,
$\sum_{(i,j)\in\mathcal{P}(B,s_{l,m}(B))}\sqrt{2^{m}(s^{m}_{i,j+1}(B)-s^{m}_{i,j}(B))}$ $\leq c_{6}\log(T)\sqrt{2^{m}(s_{i,j_{0}}-\tau_{l,m})}$
$\leq c_{6}\log(T)4^{m}\,.$
Since Proposition 10 holds uniformly over all bins $B\in\mathcal{T}_{m}$, it also holds for the particular choice $B=B^{\rm{last}}_{l,m}$, and hence in both cases $s_{l,m}(B)=\tau_{l,m}+8^{m}$ and $s_{l,m}(B)<\tau_{l,m}+8^{m}$, we have
$\sum_{(i,j)\in\mathcal{P}(B^{\rm{last}}_{l,m},s_{l,m}(B^{\rm{last}}_{l,m}))}\sqrt{2^{m}(s^{m}_{i,j+1}(B^{\rm{last}}_{l,m})-s^{m}_{i,j}(B^{\rm{last}}_{l,m}))}\leq c_{6}\log(T)4^{m}\,.$
Now recall by Definition 12 that $s_{i,j_{0}+1}^{m}(B^{\rm{last}}_{l,m})$ is the earliest round satisfying the lower bound on the cumulative regret
$\sum_{t=s^{m}_{i,j_{0}}(B^{\rm{last}}_{l,m})}^{s^{m}_{i,j_{0}+1}(B^{\rm{last}}_{l,m})}\delta_{t}(B_{t,m}^{\sharp},B^{\rm{last}}_{l,m})\geq c_{5}\log(T)\sqrt{2^{m}\left(s^{m}_{i,j_{0}}(B^{\rm{last}}_{l,m})-s^{m}_{i,j_{0}-1}(B^{\rm{last}}_{l,m})\right)}\,,$
@@
$\sum_{t=s^{m}_{i,j_{0}}(B^{\rm{last}}_{l,m})}^{s^{m}_{i,j_{0}+1}(B^{\rm{last}}_{l,m})-1}\delta_{t}(B_{t,m}^{\sharp},B^{\rm{last}}_{l,m})\leq c_{5}\log(T)\sqrt{2^{m}\left(s^{m}_{i,j_{0}}(B^{\rm{last}}_{l,m})-s^{m}_{i,j_{0}-1}(B^{\rm{last}}_{l,m})-1\right)}\,.$
@@
$\sum_{(i,j)\in\mathcal{P}(B^{\rm{last}}_{l,m},s_{l,m}(B^{\rm{last}}_{l,m}))}\sum_{t=s_{i,j}^{m}(B^{\rm{last}}_{l,m})}^{s_{i,j+1}^{m}(B^{\rm{last}}_{l,m})}\delta_{t}(B_{t,m}^{\sharp},B^{\rm{last}}_{l,m})\leq c_{5}c_{6}\log^{2}(T)4^{m}\,.$ (20)
It remains to bound the regret incurred over the non-bad segments. To this end, we observe from Definition 12 that each phase contains at most one such segment. Therefore, there exists a positive numerical constant $c_{7}$ such that
$\sum_{I\subseteq[\tau_{l,m},\tau_{l,m+1}[\text{ s.t. }I\text{ is non-bad segment}}\,\sum_{t\in I}\,\delta_{t}(B_{t,m}^{\sharp},B^{\rm{last}}_{l,m})\leq c_{7}\log(T)4^{m}\,.$ (21)
Summing Equations 20 and 21 gives the total relative regret of $B^{\rm{last}}_{l,m}$ with respect to $B_{t,m}^{\sharp}$ over the block on event $\mathcal{E}_{1}\cap\mathcal{E}_{2}(\tau_{l,m})$,
$\sum_{t=\tau_{l,m}}^{\tau_{l,m+1}-1}\delta_{t}(B_{t,m}^{\sharp},B^{\rm{last}}_{l,m})\leq\overbrace{\left(c_{5}c_{6}+c_{7}\right)}^{{\color[rgb]{0.68,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.68,0,0}\pgfsys@color@cmyk@stroke{0}{0.87}{0.68}{0.32}\pgfsys@color@cmyk@fill{0}{0.87}{0.68}{0.32}=c_{8}}}\log^{2}(T)4^{m}\,.$
Finally, we use the fact that event $\mathcal{E}_{2}(\tau_{l,m})$ holds with probability at least $1-1/T^{2}$ (Proposition 10),
@@
Choosing ${\color[rgb]{0.68,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.68,0,0}\pgfsys@color@cmyk@stroke{0}{0.87}{0.68}{0.32}\pgfsys@color@cmyk@fill{0}{0.87}{0.68}{0.32}c_{B}=c_{8}+1}$ concludes the proof.
@@ D.8 Conclusion of the proof
Recall from Section D.5 that we have
$\mathbb{E}\left[R(\pi,T)\right]$ $\leq\log(T)\sum_{i=0}^{\tilde{L}_{T}}(\tau_{i+1}-\tau_{i})^{2/3}+\frac{7}{T}$
@@
Propositions 7 and 8 prove that the last sum is upper bounded as
@@
where we used Equation 10 for the second inequality, and where we reconditioned on event $\mathcal{E}_{1}$ in the last line. Therefore,
$\mathbb{E}\left[R(\pi,T)\right]$ $\leq\log(T)\sum_{i=0}^{\tilde{L}_{T}}(\tau_{i+1}-\tau_{i})^{2/3}+\frac{7+c_{A}+c_{B}}{T}$
@@
We conclude the proof by relating episodes and phases, which would permit to bound the term
@@
We define the phases intersecting episode $[t_{l},t_{l+1}[$ as
@@
Therefore, we can rewrite the upper bound on cumulative regret over each episode as
@@
Recall that Proposition 6 shows that on $\mathcal{E}_{1}$, each phase $[\tau_{i},\tau_{i+1}[$ intersects at most two episodes. Therefore,
@@
$\mathbb{E}\left[R(\pi,T)\right]$ $\leq\log(T)\sum_{i=0}^{\tilde{L}_{T}}(\tau_{i+1}-\tau_{i})^{2/3}+\frac{7+c_{A}+c_{B}}{T}$
$+12\sum_{i=1}^{\tilde{L}_{T}}(\tau_{i+1}-\tau_{i})^{2/3}+12(c_{A}+c_{B})\log^{2}(T)\sum_{i=1}^{\tilde{L}_{T}}(\tau_{i+1}-\tau_{i})^{2/3}\,.$
Appendix E Proof of Corollary 1
@@
Fix a phase $[\tau_{i},\tau_{i+1}[$, where $\tau_{i+1}\leq T+1$. Define the total variation within this phase as
@@
Let $x_{i}=\operatorname*{argmax}_{x\in[0,1]}\mu_{\tau_{i+1}}(x)$ denote the best arm at the end of phase $i$. By definition, there exists an interval $[s_{1},s_{2}]\subset[\tau_{i},\tau_{i+1}[$ over which $x_{i}$ suffers significant regret:
@@
We now recall the upper bound from Theorem 2:
$\mathbb{E}\left[R(\pi,T)\right]\leq\bar{c}\log^{2}(T)\sum_{i=0}^{\tilde{L}_{T}}(\tau_{i+1}-\tau_{i})^{2/3}\,.$
@@
This completes the proof.
Appendix F Proof of the lower bound (Theorem 1)
Proof sketch. The proof of Theorem 1 relies on the following ideas. First, note that the lower bound $T^{2/3}$ follows from classical results (Kleinberg, 2004). To prove the lower bound either with respect to total variation $V_{T}$, or with respect to the total number of variations $L_{T}$, we construct $T/\tau$ stationary sub-problems with horizon $\tau$ to be chosen later. For each problem, we design $K=\tau^{1/3}$ possible Lipschitz-continuous rewards function. Each of these functions is almost flat, with the exception of one bump of size $\epsilon\approx\tau^{-1/3}$ hidden among $K$ possible bins. Classical arguments show that any bandit algorithm will misidentify the optimal bin with constant probability in some problem instance, thus suffering a regret $\tau^{2/3}$. Then, we create a dynamic problem instance by concatenating $T/\tau$ such problems, where the reward function corresponding to each problem of horizon $\tau$ is chosen independently. The total regret of any algorithm must then be $T/\tau\times\tau^{2/3}=T\tau^{-1/3}$.
@@ Appendix F Proof of the lower bound (Theorem 1)
Sequential construction of the non-stationary reward function. For a given algorithm $\pi$, we build a difficult non-stationary reward function sequentially. We show by induction that for all $l\leq L$, there exists a non-stationary sequence of reward functions $(\mu_{t})_{t\leq l\tau}$ such that the expected cumulatiive regret of algorithm $\pi$ over the stationary phases $\mathcal{P}_{1}$, …, $\mathcal{P}_{l}$ is at least $0.01K^{2}l$ for all $l\in[\![L]\!]$. We abuse notation and define $\mathcal{P}_{0}=\emptyset$, so that the base case of the recursion (corresponding to $l=0$) becomes trivial.
@@
Now, if $k^{*}=1$, it means that under $\mu^{1}$, the algorithm $\pi$ chooses in expectation at least $\frac{(K-1)\tau}{K}$ actions outside of interval $I_{1}$ in phase $l+1$. Then, the regret in this phase must be at least $\frac{(K-1)\tau}{K}\times\frac{\epsilon}{4}\geq\frac{\tau\epsilon}{8}=K^{2}/16$. In this case, using the assumption that the expected cumulated regret of algorithm $\pi$ over the stationary phases $\mathcal{P}_{1}$, …, $\mathcal{P}_{l}$ is at least $0.01K^{2}l$, we find that the cumulative rewards up to phase $\mathcal{P}_{l+1}$ is at least $0.01K^{2}(l+1)$, thus completing the induction step. We therefore assume henceforth that $k^{*}\neq 1$. Now, denoting $\mathcal{K}\left(\mathbb{P},\mathbb{Q}\right)$ the Kullback-Leibler divergence between probabilities $\mathbb{P}$ and $\mathbb{Q}$, and by $kl(p,q)$ the Kullback-Leibler divergence between Bernoulli distributions with parameters $p$ and $q$, classical arguments (see, e.g., (Lattimore and Szepesvári, 2020, Chapter 15.1)) show that
@@
where we have used the definition of $k^{*}$ and of $\tau$ and $\epsilon$. Now, we apply Bretagnolle-Huber Inequality (see, e.g., (Lattimore and Szepesvári, 2020, Chapter 14)) to the $\mathcal{F}_{\tau(l+1)}$-measurable event $A=\{N_{1}^{l+1}\leq\frac{\tau}{2}\}$. This yields
@@
Lower bound depending on $L_{T}$. For $L_{T}\in\{0,1,2\}$, the bound follows from classical results on Lipschitz-continuous bandits (see, e.g., Kleinberg (2004)), by noticing that since there exists a stationary problem and a constant $c>0$ on which any bandit algorithm must have regret at least $cT^{2/3}$, there exists a problem with 0 shifts (i.e., at most $L_{T}$) such that the regret is at least $cL^{1/3}T^{2/3}/2$. Therefore, we assume that $L_{T}>2$. Then, setting $K=\lfloor(T/L_{T})^{1/3}\rfloor$, we see that $K^{3}\leq T/2$. Thus, the non-stationary reward function designed above has regret at least $0.005T/K\geq 0.005T^{2/3}L_{T}^{1/3}$. Moreover, this reward function has at most $L_{T}$ changes.
Lower bound depending on $V_{T}$. For $V_{T}\leq 2T^{-1/3}$, we have $V_{T}^{1/4}T^{3/4}\leq 2T^{2/3}$, the result follows from classical results on Lipschitz-continuous bandits (see, e.g., Kleinberg (2004)) by noticing that there exists a stationary environment and a constant $c>0$ on which any bandit algorithm must have regret at least $cT^{2/3}$. Therefore, we assume henceforth that $V_{T}>T^{-1/3}$. Then, setting $K=\lfloor(T/V_{T})^{1/4}\rfloor$, we see that $K^{3}\leq T/2$. Thus, the non-stationary reward function designed above has regret at least $0.005T/K\geq 0.005T^{3/4}V_{T}^{1/4}$. Moreover, the cumulative variation of this reward function is at most $V_{T}$.
Appendix G Extension to multi-dimensional action space
In this section, we discuss how to extend our algorithm and results to multi-dimensional spaces.
Problem setting in dimension $p$. We assume that at each round, the player chooses an action in the set $[0,1]^{p}$, where $p\in\mathbb{N}_{*}$ is the dimension. We assume that each mean reward function $\mu_{t}:[0,1]^{p}\to[0,1]$ satisfies the following Lipschitz condition:

### v2

@@ D.7 Upper bounding (B)
Proof of Proposition˜10.
Let $\tau_{l,m},\tau_{l,m+1}[$ be a block, and let $B\in\mathcal{T}_{m}$ such that $s_{l,m}(B)<\tau_{l,m}+8^{m}$. Let $d_{i,j}$ the integer satisfying
@@
$\geq\frac{c_{4}}{8}\log(T)\,$,
where we use the fact that there is at least one round $t\in[s_{i,j}^{m}(B),\tilde{s}_{i,j}^{m}(B)]$ such that $t-\tau_{l,m}\equiv 0[8^{d_{ij}}]$, i.e. that at least a replay at depth $d_{i,j}$ that can be scheduled in $s_{i,j}^{m}(B),s_{i,j+1}(B)]$, and where we used Definition˜13 in the last inequality. Setting $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}c_{4}=192$}, we have
@@
Therefore, using Proposition˜10, for any bin $B\in\mathcal{T}_{m}$, on event $\mathcal{E}_{1}\cap\mathcal{E}_{2}(\tau_{l,m})$, we have that there exists a bad segment $s_{i,j_{0}}^{m}(B),s_{i,j_{0}+1}^{m}(B)]$ such that $s_{i,j_{0}+1}^{m}(B)<s_{l,m}(B)<\tau_{l,m+1}$, and there exists a $\texttt{Replay}(t_{\mathrm{start}},d_{i,j_{0}})$ starting within this segment, $t_{\mathrm{start}}\in[s_{i,j_{0}}^{m}(B),s_{i,j_{0}+1}^{m}(B)]$ and $d_{i,j_{0}}$ satisfying $8^{d_{i,j_{0}}}\leq s_{i,j_{0}+1}^{m}(B)-s_{i,j_{0}}^{m}(B)\leq 8^{d_{i,j_{0}}+2}$. By Proposition˜9, this implies that bin $B$ is evicted by round $s_{i,j_{0}+1}^{m}(B)$. Therefore, on $\mathcal{E}_{1}\cap\mathcal{E}_{2}(\tau_{l,m})$, using the definition of the bad round $s_{l,m}(B)<\tau_{l,m+1}$,
$\sum_{(i,j)\in\mathcal{P}(B,s_{l,m}(B))}\sqrt{2^{m}(s^{m}_{i,j+1}(B)-s^{m}_{i,j}(B))}$ $\leq c_{4}\log(T)\sqrt{2^{m}(s_{i,j_{0}}-\tau_{l,m})}$
$\leq c_{4}\log(T)4^{m}\,.$
Since Proposition˜10 holds uniformly over all bins $B\in\mathcal{T}_{m}$, it also holds for the particular choice $B=B^{\rm{last}}_{l,m}$, and hence in both cases $s_{l,m}(B)=\tau_{l,m}+8^{m}$ and $s_{l,m}(B)<\tau_{l,m}+8^{m}$, we have
$\sum_{(i,j)\in\mathcal{P}(B^{\rm{last}}_{l,m},s_{l,m}(B^{\rm{last}}_{l,m}))}\sqrt{2^{m}(s^{m}_{i,j+1}(B^{\rm{last}}_{l,m})-s^{m}_{i,j}(B^{\rm{last}}_{l,m}))}\leq c_{4}\log(T)4^{m}\,.$
Now recall by Definition˜12 that $s_{i,j_{0}+1}^{m}(B^{\rm{last}}_{l,m})$ is the earliest round satisfying the lower bound on the cumulative regret
$\sum_{t=s^{m}_{i,j_{0}}(B^{\rm{last}}_{l,m})}^{s^{m}_{i,j_{0}+1}(B^{\rm{last}}_{l,m})}\delta_{t}(B_{t,m}^{\sharp},B^{\rm{last}}_{l,m})\geq c_{3}\log(T)\sqrt{2^{m}\left(s^{m}_{i,j_{0}}(B^{\rm{last}}_{l,m})-s^{m}_{i,j_{0}-1}(B^{\rm{last}}_{l,m})\right)}\,$,
@@
$\sum_{t=s^{m}_{i,j_{0}}(B^{\rm{last}}_{l,m})}^{s^{m}_{i,j_{0}+1}(B^{\rm{last}}_{l,m})-1}\delta_{t}(B_{t,m}^{\sharp},B^{\rm{last}}_{l,m})\leq c_{3}\log(T)\sqrt{2^{m}\left(s^{m}_{i,j_{0}}(B^{\rm{last}}_{l,m})-s^{m}_{i,j_{0}-1}(B^{\rm{last}}_{l,m})-1\right)}\,.$
@@
$\sum_{(i,j)\in\mathcal{P}(B^{\rm{last}}_{l,m},s_{l,m}(B^{\rm{last}}_{l,m}))}\sum_{t=s_{i,j}^{m}(B^{\rm{last}}_{l,m})}^{s_{i,j+1}^{m}(B^{\rm{last}}_{l,m})}\delta_{t}(B_{t,m}^{\sharp},B^{\rm{last}}_{l,m})\leq c_{3}c_{4}\log^{2}(T)4^{m}\,.$ (20)
It remains to bound the regret incurred over the non-bad segments. To this end, we observe from Definition˜12 that each phase contains at most one such segment. Therefore, there exists a positive numerical constant $c_{5}$ such that
$\sum_{I\subseteq[\tau_{l,m},\tau_{l,m+1}[\text{ s.t. }I\text{ is non-bad segment}}\,\sum_{t\in I}\,\delta_{t}(B_{t,m}^{\sharp},B^{\rm{last}}_{l,m})\leq c_{5}\log(T)4^{m}\,.$ (21)
Summing Equations˜20 and 21 gives the total relative regret of $B^{\rm{last}}_{l,m}$ with respect to $B_{t,m}^{\sharp}$ over the block on event $\mathcal{E}_{1}\cap\mathcal{E}_{2}(\tau_{l,m})$,
$\sum_{t=\tau_{l,m}}^{\tau_{l,m+1}-1}\delta_{t}(B_{t,m}^{\sharp},B^{\rm{last}}_{l,m})\leq\overbrace{\left(c_{3}c_{4}+c_{5}\right)}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}=c_{6}}}\log^{2}(T)4^{m}\,.$
Finally, we use the fact that event $\mathcal{E}_{2}(\tau_{l,m})$ holds with probability at least $1-1/T^{2}$ (Proposition˜10),
@@
Choosing $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}c_{B}=c_{6}+1$} concludes the proof.
@@ D.8 Conclusion of the proof
Recall from Section˜D.5 that we have
$\mathbb{E}\left[R(\pi_{\texttt{MDBE}},T)\right]$ $\leq\log(T)\sum_{i=0}^{\tilde{L}_{T}}(\tau_{i+1}-\tau_{i})^{2/3}+\frac{7}{T}$
@@
Propositions˜7 and 8 prove that the last sum is upper bounded as
@@
where we used Equation˜10 for the second inequality, and where we reconditioned on event $\mathcal{E}_{1}$ in the last line. Therefore,
$\mathbb{E}\left[R(\pi_{\texttt{MDBE}},T)\right]$ $\leq\log(T)\sum_{i=0}^{\tilde{L}_{T}}(\tau_{i+1}-\tau_{i})^{2/3}+\frac{7+c_{A}+c_{B}}{T}$
@@
We conclude the proof by relating episodes and phases, which will permit to bound the final term
@@
We define the phases intersecting episode $t_{l},t_{l+1}[$ as
@@
With this definition, we can rewrite the upper bound on cumulative regret over each episode as
@@
Recall that Proposition˜6 shows that on $\mathcal{E}_{1}$, each phase $\tau_{i},\tau_{i+1}[$ intersects at most two episodes. Therefore,
@@
$\mathbb{E}\left[R(\pi_{\texttt{MDBE}},T)\right]$ $\leq\log(T)\sum_{i=0}^{\tilde{L}_{T}}(\tau_{i+1}-\tau_{i})^{2/3}+\frac{7+c_{A}+c_{B}}{T}$
$+12\sum_{i=1}^{\tilde{L}_{T}}(\tau_{i+1}-\tau_{i})^{2/3}+12(c_{A}+c_{B})\log^{2}(T)\sum_{i=1}^{\tilde{L}_{T}}(\tau_{i+1}-\tau_{i})^{2/3}$
$\leq\left(\log(T)+12+12(c_{A}+c_{B})\log^{2}(T)\right)\sum_{i=1}^{\tilde{L}_{T}}(\tau_{i+1}-\tau_{i})^{2/3}+\frac{7+c_{A}+c_{B}}{T}$
$\leq\bar{c_{1}}\log^{2}(T)\sum_{i=1}^{\tilde{L}_{T}}(\tau_{i+1}-\tau_{i})^{2/3}\,$,
where $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\bar{c_{1}}=12(c_{A}+c_{B})+2$}. This completes the proof. We summarize in Table˜2 the exact values of the numerical constants used in the proof.
Constant Numerical value
$c_{0}$ $3+7(e-1)\sqrt{2}$
$c_{1}$ $7(e-1)\sqrt{2}$
$c_{2}$ $c_{0}+c_{1}+8$
$c_{3}$ $140+4c_{1}$
$c_{4}$ 192192
$c_{5}$ Equation˜21
$c_{6}$ $c_{3}c_{4}+c_{5}$
$c_{A}$ $18(c_{0}+c_{1}+10)$
$c_{B}$ $c_{6}+1$
$\bar{c_{1}}$ $12(c_{A}+c_{B})+2$
Table 2: Summary of numerical constants used in the proof
Appendix E Proof of Corollary˜1
@@
Fix a phase $\tau_{i},\tau_{i+1}[$, where $\tau_{i+1}\leq T+1$. Define the total variation within this phase as
@@
Let $x_{i}=\operatorname*{argmax}_{x\in[0,1]}\mu_{\tau_{i+1}}(x)$ denote the best arm at the end of phase $i$. By definition, there exists an interval $s_{1},s_{2}]\subset[\tau_{i},\tau_{i+1}[$ over which $x_{i}$ suffers significant regret:
@@
We now recall the upper bound from Theorem˜2:
$\mathbb{E}\left[R(\pi_{\texttt{MDBE}},T)\right]\leq\bar{c}\log^{2}(T)\sum_{i=0}^{\tilde{L}_{T}}(\tau_{i+1}-\tau_{i})^{2/3}\,.$
@@
Appendix F Proof of the lower bound (Theorem˜1)
Proof sketch. The proof of Theorem˜1 relies on the following ideas. First, note that the lower bound $T^{2/3}$ follows from classical results of Lipschitz bandits [Kleinberg, 2004]. To prove the lower bound either with respect to total variation $V_{T}$, or with respect to the total number of changes $L_{T}$, we construct $T/\tau$ stationary sub-problems with horizon $\tau$ to be chosen later. For each problem, we design $K=\tau^{1/3}$ possible Lipschitz-continuous rewards function. Each of these functions is almost flat, with the exception of one bump of size $\epsilon\approx\tau^{-1/3}$ hidden among $K$ possible bins. Classical arguments show that any bandit algorithm will misidentify the optimal bin with constant probability in some problem instance, thus suffering a regret $\tau^{2/3}$. Then, we create a dynamic problem instance by concatenating $T/\tau$ such problems, where the reward function corresponding to each problem of horizon $\tau$ is chosen independently. The total regret of any algorithm must then be
$\frac{T}{\tau}\tau^{2/3}=T\tau^{-1/3}\,.$
@@ Appendix F Proof of the lower bound (Theorem˜1)
Sequential construction of the non-stationary reward function. For a given algorithm $\pi$, we build a difficult non-stationary reward function sequentially. We show by induction that for all $l\leq L$, there exists a non-stationary sequence of reward functions $(\mu_{t})_{t\leq l\tau}$ such that the expected cumulative regret of algorithm $\pi$ over the stationary phases $\mathcal{P}_{1}$, …, $\mathcal{P}_{l}$ is at least $0.01K^{2}l$ for all $l\in[\![L]\!]$. We abuse notation and define $\mathcal{P}_{0}=\emptyset$, so that the base case of the recursion (corresponding to $l=0$) becomes trivial.
@@
Now, if $k^{*}=1$, it means that under $\mu^{1}$, the algorithm $\pi$ chooses in expectation at least $\frac{(K-1)\tau}{K}$ actions outside of interval $I_{1}$ in phase $l+1$. Then, the regret in this phase must be at least $\frac{(K-1)\tau}{K}\times\frac{\epsilon}{4}\geq\frac{\tau\epsilon}{8}=K^{2}/16$. In this case, using the assumption that the expected cumulated regret of algorithm $\pi$ over the stationary phases $\mathcal{P}_{1}$, …, $\mathcal{P}_{l}$ is at least $0.01K^{2}l$, we find that the cumulative rewards up to phase $\mathcal{P}_{l+1}$ is at least $0.01K^{2}(l+1)$, thus completing the induction step. We therefore assume henceforth that $k^{*}\neq 1$. Now, denoting $\mathcal{K}\left(\mathbb{P},\mathbb{Q}\right)$ the Kullback-Leibler divergence between probabilities $\mathbb{P}$ and $\mathbb{Q}$, and by $kl(p,q)$ the Kullback-Leibler divergence between Bernoulli distributions with parameters $p$ and $q$, classical arguments (see, e.g., [Lattimore and Szepesvári, 2020, Chapter 15.1]) show that
@@
where we have used the definition of $k^{*}$ and of $\tau$ and $\epsilon$. Now, we apply Bretagnolle-Huber Inequality (see, e.g., [Lattimore and Szepesvári, 2020, Chapter 14]) to the $\mathcal{F}_{\tau(l+1)}$-measurable event $A=\{N_{1}^{l+1}\leq\frac{\tau}{2}\}$. This yields
@@
Lower bound depending on $L_{T}$. For $L_{T}\in\{0,1,2\}$, the bound follows from classical results on Lipschitz-continuous bandits (see, e.g., Kleinberg [2004]), by noticing that since there exists a stationary problem and a constant $c>0$ on which any bandit algorithm must have regret at least $cT^{2/3}$, there exists a problem with 0 shifts (i.e., at most $L_{T}$) such that the regret is at least $cL^{1/3}T^{2/3}/2$. Therefore, we assume that $L_{T}>2$. Then, setting $K=\lfloor(T/L_{T})^{1/3}\rfloor$, we see that $K^{3}\leq T/2$. Thus, the non-stationary reward function designed above has regret at least $0.005T/K\geq 0.005T^{2/3}L_{T}^{1/3}$. Moreover, this reward function has at most $L_{T}$ changes.
Lower bound depending on $V_{T}$. For $V_{T}\leq 2T^{-1/3}$, we have $V_{T}^{1/4}T^{3/4}\leq 2T^{2/3}$, the result follows from classical results on Lipschitz-continuous bandits (see, e.g., Kleinberg [2004]) by noticing that there exists a stationary environment and a constant $c>0$ on which any bandit algorithm must have regret at least $cT^{2/3}$. Therefore, we assume henceforth that $V_{T}>T^{-1/3}$. Then, setting $K=\lfloor(T/V_{T})^{1/4}\rfloor$, we see that $K^{3}\leq T/2$. Thus, the non-stationary reward function designed above has regret at least $0.005T/K\geq 0.005T^{3/4}V_{T}^{1/4}$. Moreover, the cumulative variation of this reward function is at most $V_{T}$.
Appendix G Extensions
G.1 Extension to multi-dimensional action space
In this subsection, we discuss how to extend our algorithm and results to multi-dimensional spaces.
Problem setting in dimension $p$. We assume that at each round, the player chooses an action in the set $0,1]^{p}$, where $p\in\mathbb{N}_{*}$ is the dimension. We assume that each mean reward function $\mu_{t}:[0,1]^{p}\to[0,1]$ satisfies the following Lipschitz condition: